A community water tank can be filled in 10 hours by the town office well alone and in 16 hours by the high school well alone. How long will it take to fill the tank if both wells are working together?
Answer by nerdybill(5411)

A community water tank can be filled in 10 hours by the town office well alone and in 16 hours by the high school well alone. How long will it take to fill the tank if both wells are working together?
.
Rate of town office well: 1 tank per 10 hours : 1/10
Rate of high school well: 1 tank per 16 hours : 1/16
.
Let x = number of hours required for both wells
then
x(1/10 + 1/16) = 1
x/10 + x/16) = 1
Multiplying both sides by 160 (common denominator):
16x + 10x = 160
26x = 160
x = 160/26
x = 6.154 hours